OpenStudy (credmond):

Find the point on the graph of f(x)= sqrt(x) where the tangent line has a slopw of 1/8. The answer is (16, 4) but I am not sure how to get to it. Can someone help me please!? :)

OpenStudy (anonymous):

st f'(x) = 1/8

OpenStudy (anonymous):

set*

OpenStudy (credmond):

So do you do 1/8 = x ^1/2?

OpenStudy (anonymous):

no f'(x) = (1/2) x^(-1/2)

OpenStudy (credmond):

ohhh sorry I didn't see the ' symbol

OpenStudy (credmond):

So what do you do after you have the derivative?

OpenStudy (credmond):

Oh, never mind I figured it out! Thank you!